### v2

@@ 3. Proof of Theorem 9
First of all, we suppose, without loss of generality, that has finite order Combining this with Lemma 1, Lemma 2 (i) and the assumption that and are two distinct non-constant meromorphic functions that share IM and CM, we deduce that and are two distinct transcendental meromorphic functions in the complex plane.
From Lemma 2(i), Remark 4 and the assumption we have
@@
By (17) and Whittaker [41, p. 82] we have
By (18), (19) and Lemma 4(i) we have
@@
We consider the following two cases:
@@
On the other hand, by (17), Lemma 4 (i) and the second fundamental theorem (cf.[12, Theorem 2.1]) we have
Since and are transcendental meromorphic functions, by (29) and 5 we deduce that at least one of and say, satisfies
According to (30), Lemma 3 (ii) and the assumption that and share IM and CM, we have a discussion as follows:
@@
where and in what follows, denotes the reduced counting function of the common simple zeros of and in
By (31) we see that there exists
an a infinite sequence of common simple zeros of and in such that as Therefore, it follows by (28) that for each Combining this with the fact that is a non-vanishing rational function, we deduce that is reduced to the constant and so (28) can be rewritten into
By (32) and the assumption that and share IM we deduce that and share CM. This together with the assumption that and share CM, we get the assertion of Theorem 9.
@@
where and in what follows, denotes the reduced counting function of those common zeros of and in and each such point in is a simple zero of and a zero of of multiplicity By (33) we see that there exists an infinite sequence of those common zeros of and in such that as Here each such point is a simple zero of and a zero of of multiplicity Combining this with (28), we deduce
Since is a non-vanishing rational function, we deduce by (34) that is reduced to the constant and so (28) can be rewritten into
From (35), the assumption that and share IM, and the supposition that is a positive integer such that we deduce that if is a zero of of multiplicity with and then is a zero of of multiplicity with and This implies that the functions
@@
are two entire functions. Therefore, from (17), (36) and Lemma 4 (i), we deduce
@@
From (35) and the assumption that and share IM, we deduce that if with and is a common zero of and with and then is a zero of of multiplicity larger than or equal to with for and Therefore, from (17), Lemma 2(ii), Remark 2, Lemma 4 (i) and the first fundamental theorem, we deduce
@@
From (42), (45) and the Valiron-Mokhonko lemma (cf.[23]) we deduce
@@
We take the Nevanlinnas characteristic functions on both sides of (55), and then use the Valiron-Mokhonko lemma (cf.[23]) to get
@@
We take the Nevanlinnas characteristic functions on both sides of (57) and then use the Valiron-Mokhonko lemma (cf.[23]) to get (56).
@@
We take the Nevanlinnas characteristic functions on both sides of (58) and then use the supposition and the Valiron-Mokhonko lemma (cf.[23]) to get
@@
where and in what follows, denotes the reduced counting function of those common zeros of and in and each such point in is a zero of of multiplicity and a simple zero of By (60) we see that there exists an infinite sequence of those common zeros of and in such that as Here each such point is a zero of of multiplicity and a simple zero of Combining this with (28), we deduce
Since is a non-vanishing rational function, and is a complex constant, we deduce by (61) that is reduced to the constant and so (28) can be rewritten into
From (62), the assumption that and share IM, and the supposition that is a positive integer such that we deduce that if is a zero of of multiplicity with and then is a zero of of multiplicity with and This implies that the functions
@@
are two entire functions. Next we use (17), (63), (64), Lemma 4 (i), the lines of (38) and (39) in Subcase 1.2 to deduce that either with or that with is reduced to a non-vanishing polynomial. Suppose that one of and holds. Then, we use (62)-(64) and the similar reasonings as in Subcase 1.2 to get a contradiction. Next we suppose that and are non-vanishing polynomials. Then, we use (63), (64), the the similar reasonings from the line before (41) to the line of (45) to deduce
@@
From (62) and the assumption that and share IM, we deduce that if with and is a common zero of and with and then is a zero of of multiplicity larger than or equal to with for and Therefore, from (17), Lemma 2(ii), Remark 2, Lemma 4 (i) and the first fundamental theorem, we deduce
@@
By (65), (66) and the Valiron-Mokhonko lemma (cf.[23]) we deduce
@@
where with are finite complex constants, and is a positive integer. Next we let with and and consider the following angles for any given positive number satisfying
@@
for
On the other hand, since and are non-constant meromorphic functions of finite order such that they share IM and CM, we have by (17), Lemma 3(ii) and Remark 4 that
We consider the following two subcases:
@@
for the the even integer satisfying and the positive number such that From (17) and Remark 5 we have
@@
Next we use Lemma 9 and the reasonings in the proof of Lemma 2.2.2 in [49, pp.53-54] to deduce
for the even integer satisfying and the positive number such that Here and in what follows, for the integer satisfying while is defined as with being the number of zeros of for in counted with multiplicities. From (91) and (92) we have
@@
Now we use (86) and (87) to deduce for the positive integer satisfying Combining this with (83), we deduce that for any given positive number satisfying there exists some positive integer with that satisfies
@@
for the positive integer satisfying with From (83), (97) and the known equality with the even integer satisfying we have
@@
From (17) and Remark 5 we have
@@
[49, pp.53-54, Lemma 2.2.2] to deduce
for the odd integer satisfying and the positive number such that Here and in what follows, for the integer satisfying while is defined as with being the number of zeros of for in counted with multiplicities. From (103) and (104) we have
@@
for the odd integer satisfying and the positive integer satisfying with From (83), (107) and the known equality with the odd integer satisfying we deduce (98) for the odd integer satisfying and the given positive number satisfying From (98) we get a contradiction. This proves Claim 2I.
Last, we use Lemma 12, Lemma 13 and Claim 2I to complete the proof of Theorem 9. For this purpose, we have a discussion as follows: suppose that Then, it follows from Lemma 2.2 in [13] that there exist some real constant and some set that has infinite linear measure such that for all Combining this with Theorem 7, we see that and share all four values
@@
From (108) we see that there exists some positive number satisfying
such that
where and in what follows, is a positive constant such that
@@
for each integer satisfying From (109)-(111) we have
@@
Next we let
and let be the positive boundary of the simply connected domain where and are two oriented segments that are defined as
@@
respectively, and is a directional arc that is defined as
@@
where and in what follows, and are defined as in (82) for
Next we define as
Here and in what follows, is a small positive number satisfying such that Noting that with is continuous uniformly for on the closed interval and continuous at the origin point we deduce from (110) and (111) with (113) and (118) that
for the point the small positive number satisfying the positive number satisfying and the large positive number satisfying Here and in what follows, is defined as in (110) for From (110) and (111) with (113)-(119), the definitions of the proximity function and the Nevanlinna characteristic function of with the center at for we have
for the small positive number satisfying the positive number satisfying and the large positive number satisfying On the other hand, from Lemma 12, Claim 2I and the definition of the counting function of with the center at for we have
@@
with and for and From (122), Claim 2I and Lemma 13 we see that there exists a set
that has linear measure zero, such that if then there is a constant, say that depend only upon and such that for all satisfying and and for each we have
From the formula (2.1.1) in [49, p.27], we have
@@
for with all satisfying
From (122) and (125) we deduce
for with all satisfying
From (121) and (126) we have
for with all satisfying
From (120) and (127) we have
for with all satisfying
From (128), Lemma 5, the supposition that is a positive number satisfying and the known result that and are transcendental meromorphic functions, we get a contradiction. This completes the proof of Theorem 9.

### v7

@@ 3. Proof of Theorem 7
First, before the formal beginning of the proof of Theorem 7, we want to point out that the lines from the next formal beginning of the proof of Theorem 7 to the end of Case 1 of the next formal beginning of the proof of Theorem 7 are covered essentially by the proof of Steinmetz [46, Theorem 6.2] even without any assumption on the order of growth. For the convenience of the readers, we give the detailed proof for this part on the assumption of finite order of growth as follows: first of all, we suppose, without loss of generality, that has finite order Combining this with Lemma 1, Lemma 2 (i) and the assumption that and are two distinct non-constant meromorphic functions that share IM and CM, we deduce that and are two distinct transcendental meromorphic functions in the complex plane. From Lemma 2(i), Remark 4 and the assumption we have
@@
By (20) and Whittaker [50, p. 82] we have
By (21), (22) and Lemma 4 we have
@@
On the other hand, from (20), Lemma 2(i), Lemma 3(ii), Lemma 4, Remark 2-Remark 4 and the assumption that and are non-constant meromorphic functions that share IM and CM, we have
from we have
and
Here and in what follows, denotes the counting function of the multiple poles of in From (31), (32) and Lemma 6 we have
as From (30) we consider the following two cases:
@@
According to (30), (33) and the assumption that and share IM and CM, we have a discussion as follows:
@@
Here and in what follows, denotes the reduced counting function of the common simple zeros of and in for the value From (35) we suppose, without loss of generality, that
From (36) we see that there exists
an a infinite sequence of the common simple zeros of and in the complex plane, such that as Then, it follows by (34) that for each Combining this with the fact that is a non-vanishing rational function, we deduce that is reduced to the constant and so (34) can be rewritten into
From (37) and the assumption that and share IM we deduce that and share CM. This together with the assumption that and share CM, we get the assertion of Theorem 7.
@@
where and in what follows, denotes the reduced counting function of those common zeros of and in that are simple zeros of and are zeros of of multiplicity for From (38) we suppose, without loss of generality, that
From (39) we see that there exists an infinite sequence of those common zeros of and in the complex plane that are simple zeros of and are zeros of of multiplicity Combining this with (34), we deduce
Since is a non-vanishing rational function, we deduce from (40) that is reduced to the constant and so (34) can be rewritten into
From (41), the assumption that and share IM, and the supposition that is a positive integer satisfying we deduce that if is a zero of of multiplicity for and then is a zero of of multiplicity for and This implies that the functions
@@
are two entire functions. Therefore, from (20), (42) and Lemma 4, we deduce
@@
From (41) and the assumption that and share IM, we deduce that if with and is a common zero of and with and then is a zero of of multiplicity larger than or equal to with and for and Therefore, from (20), Lemma 2(ii), Remark 2, Lemma 4 and the first fundamental theorem, we deduce
@@
From (48), (51) and the Valiron-Mokhon′ko lemma (cf.[32]) we deduce
@@
We take the Nevanlinna′s characteristic functions on both sides of (61), and then use the Valiron-Mokhon′ko lemma (cf.[32]) to get
@@
We take the Nevanlinna′s characteristic functions on both sides of (63) and then use the Valiron-Mokhon′ko lemma (cf.[32]) to get (62).
@@
We take the Nevanlinna′s characteristic functions on both sides of (64) and then use the supposition and the Valiron-Mokhon′ko lemma (cf.[32]) to get
@@
where and in what follows, denotes the reduced counting function of those common zeros of and in that are zeros of of multiplicity and are simple zeros of for From (66) we suppose, without loss of generality, that
From (67) we see that there exists an infinite sequence of those common zeros of and in the complex plane
that are zeros of of multiplicity and simple zeros of Combining this with (34), we deduce
Since is a non-vanishing rational function, and is a complex constant, we deduce from (68) that is reduced to the constant and so (34) can be rewritten into
From (69), the assumption that and share IM, and the supposition that is a positive integer satisfying we deduce that if is a zero of of multiplicity with and then is a zero of of multiplicity with and This implies that the functions
@@
are two entire functions. Next we use (20), (70), (71), Lemma 4, the lines of (44) and (45) in Subcase 1.2 to deduce that either with or that with is reduced to a non-vanishing polynomial. Suppose that one of and holds. Then, we use (69)-(71) and the similar reasonings as in Subcase 1.2 to get a contradiction. Next we suppose that and are non-vanishing polynomials. Then, we use (70), (71), the the similar reasonings from the line before (47) to the line of (51) to deduce
@@
From (69) and the assumption that and share IM, we deduce that if with and is a common zero of and with and then is a zero of of multiplicity larger than or equal to with and for and Therefore, from (20), Lemma 2(ii), Remark 2, Lemma 4 and the first fundamental theorem, we deduce
@@
From (72) and (73) we deduce that if is a common pole of and then is a common zero of and Combining this with (75) and the assumption that and share CM, we deduce
By (72), (73) and the Valiron-Mokhon′ko lemma (cf.[32]) we deduce
@@
where is a positive integer, and with are finite complex constants. Next we let with and and consider the following angles for any given positive number satisfying
@@
for We consider the following two subcases:
@@
for the the even integer satisfying and the positive number such that From (20) and Lemma 12 we have
@@
Next we use Lemma 9 and the reasonings in the proof of Zheng [57, pp.53-54, Lemma 2.2.2] to deduce
for the even integer satisfying and the positive number such that Here and in what follows,
for the integer satisfying while is defined as with being the number of zeros of for in counted with multiplicities. From (97) and (98) we have
@@
Now we use (92) and (93) to deduce
for the positive integer satisfying Combining this with (90), we deduce that for any given positive number satisfying there exists some positive integer with that satisfies
@@
for the positive integer satisfying with From (90), (104) and (99) with the even integer satisfying we have
@@
From (20) and Lemma 12 we have
@@
Zheng [57, pp.53-54, Lemma 2.2.2] to have
for the odd integer satisfying and the positive number such that Here and in what follows, is defined as with being the number of zeros of for in counted with multiplicities. From (110) and (111) we have
@@
for the odd integer satisfying and the positive integer satisfying with From (90), (114) and (99) with the odd integer satisfying we deduce (105) for the odd integer satisfying and the given positive number satisfying From (105) we get a contradiction. This proves Claim 2I.
Next we use Lemma 16, Lemma 18 and Claim 2I to complete the proof of Theorem 7. For this purpose, we have a discussion as follows: suppose that Then, it follows from Lemma 2.2 in [21] that there exist some real constant and some set that has infinite linear measure such that for all Combining this with Theorem 5, we see that and share all four values
@@
for each integer satisfying From (116) and (117) we have
On the other hand, from (20), (115) and Lemma 15 we have
@@
with
for arbitrary Pó1ya
peaks of order as Here and in what follows, is some positive constant, and is some set of finite linear measure. Then, it follows from (120) and (121) that
where and in what follows, is a large positive number satisfying
Next we set the simply connected domain
with the positive boundary where and are two oriented segments such that
@@
respectively, and is a directional arc such that
@@
where and in what follows, and are defined as (89) for Next we also define with as
for Here and in what follows, is a small positive number satisfying and
is a complex number such that Then, from (123)-(128), Remark 5.3, the positive boundary of the region and the definition of the proximity function of with the center at for we have
for the small positive number satisfying Next we also set the following transformation
where and in what follows, is a simply connected domain such that
where is a positive number satisfying
and the positive boundary of the simply connected domain is
where and in what follows, is an oriented segment, and is a directional arc. They are defined as follows: the oriented segment is such an oriented segment that for each we have with and the starting point of is while the end point of is
The directional arc is such a directional arc that for each we have
with and the starting point of is with while the end point of is with In addition, from (89) with we have
From (130) and (133) we can easily verify that the transformation of (130) maps the simply connected domain of (131) onto the simply connected domain of (123) in a one-to-one mapping manner, and maps the positive boundary of into the positive boundary of in a one-to-one mapping manner,
where the oriented segment and the directional arc are mapped into and respectively by the transformation of (130) in a one-to-one mapping manner. Moreover, from (131)-(133), the polar form of the Cauchy-Riemann partial differential equations and the polar form of the sufficient conditions for an analytic function (cf.[43]), we can verify that the transformation of (130) is also an analytic transformation in the simply connected domain of (131). Combining this with the known result that the transformation of (130) maps the simply connected domain of (131) onto the simply connected domain of (123) in a one-to-one mapping manner, we see that the univalent transformation (130) on the simply connected domain is a univalent and analytic map on the simply connected domain and maps conformally the simply connected domain of (131) with the positive boundary onto the simply connected domain of (123) with the positive boundary For convinience of the following discussion, we denote
the univalent and analytic transformation (130) as
and there exists a unique point satisfying
such that the univalent and analytic mapping (134) maps conformally the simply connected domain of (131) with the positive boundary onto the simply connected domain of (123) with the positive boundary
On the other hand, from the Riemann mapping theorem (cf.[2, p.230,Theorem 1]) we also see that there exists a unique univalent and analytic function on the simply connected domain of (123), say
such that the univalent and analytic transformation (136) maps conformally the simply connected domain of (123) with the positive boundary onto the open unit disk
with the positive boundary with where the starting point of is with and the end point of is with From (134)-(136) we deduce that the composition
of the transformation (136) and the transformation (134) with for
is a univalent and analytic transformation of onto the open unit disk such that it maps the positive boundary of into the positive boundary of the open unit disk mentioned above in a one-to-one mapping manner. Noting that the transformation of (136) is a univalent and analytic mapping of onto the open unit disk Combining this with the theorem on the boundary behavior of conformal mappings, we see that the function
(137) is continuous on and for Next we use the similar reasoning of proving (8) in Remark 5 to have
and
for each point with and on the semi-circle
Since the transformation (134) is a univalent and analytic transformation of onto such that (135) holds, we have from (134), (136)-(138), (140) and Lemma 17 that the Green’s functions and for and with singularities and respectively satisfy the relation
for each
Next we use (135), (137)-(142) and the invariance of first order differential form to deduce
with for each point with and on the semi-circle
From (128), (131), (132), (141), (13) in Remark 5.1 and Goldberg-Ostrovskii [14, p.306] we have
for each point with and on the semi-circle
and the point
Here and in what follows
From (122), (129), (135), (143), (144)-(146), the supposition the definitions of the proximity function and the Nevanlinna characteristic function of with the center at for for and we deduce
i.e.,
for the small positive number satisfying and the positive number and Here and in what follows,
Next we denote by
the Nevanlinna′s angular characteristic function on the angular domain
with
Next we use (149), (150) and the line of the reasoning for proof of (97) in Subcase 2.1 of the proof of Claim 2I to deduce
for the positive number such that say. From (122), (149)-(151) and Lemma 13, we deduce that there exists some positive constant and there exists some set of finite logarithmic measure, such that when and the relation
holds uniformly in Here and in what follows
Similarly, from Lemma 13 and (97) for we deduce that there exists some positive constant and there exists some set of finite logarithmic measure, such that the relation
holds uniformly in when and Here and in what follows, is defined as (99) for From (128), (154) and the supposition we deduce
as and On the other hand, from (147), (150), (152) and (153) we have
as and Last, from Lemma 16 we have
@@
with and for Noting that denotes the Green function for the half-disc
we use the formula of the Green function (12) for the half-disc
in Remark 5.1 to have
with
for each From (130), (142) and (159) we have
for each with
From Claim 2I we see that there exist finitely many distinct zeros of with in say
such that
where is some positive integer that depends only upon the positive numbers and Next we suppose that the multiplicity of of a zero of is equal to some positive integer for and Then, from (161), (146), the reasoning in Goldberg-Ostrovskii [14, p.306], the supposition and the definition (5) of the counting function for the meromorphic function with the center at for we have
as
Here and in what follows,
for and On the other hand, from Lemma 18 we see that there exists some set
of finite logarithmic measure such that
as and From the formula (2.1.1) in Zheng [57, p.27], we have
@@
as and From (128) and (155) we deduce
as and From Remark 4, Lemma 3 (i), the assumption and the assumption that and share IM and CM we have
which implies that and have finitely many critical values including and possibly in the complex plane. Combining this with Lemma 20 and the assumption we deduce that the set of finite singular values of the inverse function is bounded. Therefore, from Lemma 19 we see that there exist constants and such that if and then
where is a positive absolute constant, in particular independent of and From (128), (155), (168) and the supposition we see that there exists some set of finite logarithmic measure such that
as and From (133) and (99) for we have Combining this with (128), (155) and (169) we deduce
as and From (128), (155), (170), Lemma 18 and the supposition
we see that there exists some set of finite logarithmic measure, such that
as and From (158), (166), (167), (171) we have
as and From (157), (158), (162), (163), (172) we have
as and as
From (146), (156) and (173) we have
as and From (148), (153), (89) and (99) for we see that
for any given small positive number satisfying
From (174), (175) and (146) we derive a contradiction, when and This completes the proof of Theorem 7.